What is the lowest value of principal quantum number(n) that allows g  orbitals to exist
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 

Quantum numbers = the set of four number which gives a complete information about the electrons in an atom.

Here n = principal quantum number

As for any value ‘n’ of principal quantum number, the Azimuthal quantum number (l) can have a value from zero to (n – 1).

For n = 1 (K shell) has l = 0 (one subshell)

For n = 2 (L shell) has l= 0,1 (2 subshell)

For n = 3 (M shell) has l = 0,1,2 (3 subshell)

For n = 4 (N shell) has l = 0,1,2,3 (4 subshell)

For n = 5 (O shell) has l = 0,1,2,3,4 (5subshell)

∴ For l = 4, minimum value of n = 5

 

 

Regards

Arun (askIITians forum expert)